A roller coaster ride in a park includes a vertical drop of 60 m. Assume that the roller coaster has a speed of nearly zero when it reaches the top of the hill. Neglecting friction, find the speed of the riders at the bottom of the hill.Β 



1
Expert's answer
2021-11-19T10:51:53-0500

We can find the speed of the riders at the bottom of the hill from the law of conservation of energy:


"KE=PE,""\dfrac{1}{2}mv^2=mgh,""v=\sqrt{2gh}=\sqrt{2\times9.8\ \dfrac{m}{s^2}\times60\ m}=34\ \dfrac{m}{s}."